dakchicken:

The two zeros of an unknown quadratic function f(x) are (-6, 0) and (8, 0) . Which of the following is the axis of symmetry of f(x) ?

Angle:

Hello @dakchicken ! The axis of symmetry is usually exactly halfway (or the average) between the x values of the given x-intercepts

dakchicken:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Angle Hello @dakchicken ! The axis of symmetry is usually exactly halfway (or the average) between the x values of the given x-intercepts \(\color{#0cbb34}{\text{End of Quote}}\) so wht is the answer?

Angle:

what is the average between -6 and 8?

dakchicken:

idk i am dumb

Angle:

\(\huge \frac{-6 + 8}{2} \) = ?

dakchicken:

-2

Angle:

incorrect what is -6 + 8 = ?

dakchicken:

2

Angle:

2 divided by 2 = ?

dakchicken:

1

Angle:

x = 1 is your answer

dakchicken:

🐔 thank you 👉👈